Filer Manufacturing has 9,284,561 shares of common stock outstanding. The current share price is $64.73, and the book value per share is $7.05. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $78,117,144, has a 0.05 coupon, matures in 10 years and sells for 83 percent of par. The second issue has a face value of $41,099,028, has a 0.06 coupon, matures in 20 years, and sells for 92 percent of par.

The most recent dividend was $1.87 and the dividend growth rate is 0.05. Assume that the overall cost of debt is the weighted average of that implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 0.34.

What is Filer's WACC? Enter the answer with 4 decimals (e.g. 0.2345)
Answer #1

Answer:

Let us calculate cost of each capital components:

Cost of equity:

Cost of equity = (Dividend next year / Price) + Dividend growth rate = (1.87 * (1 + 0.05) / 64.73) + 0.05 = 0.080333694

Cost of debt:

1st Bond:

Face value = $78,117,144

Semiannual coupon = $78,117,144 * 0.05 / 2 = $1,952,928.60

Number of semiannual periods = 10 * 2 = 20

Price = 78,117,144 * 83% = $64,837,229.52

Semiannual cost of debt = RATE (nper, pmt, pv, fv, type) = RATE (20, 1952928.60, - 64837229.52, 78117144, 0) = 3.72%

Annual cost of debt = 3.72% * 2 = 7.44%

2nd Bond:

Face value = $41,099,028

Semiannual coupon = $41,099,028 * 0.06 / 2 = 1232970.84

Number of semiannual periods = 20 * 2 = 40

Price = $41,099,028 * 92% = $37811105.76

Semiannual cost of debt = RATE (nper, pmt, pv, fv, type) = RATE (40, 1232970.84, - 37811105.76, 41099028, 0) = 3.36692261996%

Annual cost of debt = 3.36692261996% * 2 = 6.73384523992%

Weighted cost of debt = ($64,837,229.52 * 7.44% + $37811105.76 * 6.73384523992%) / ($64,837,229.52 + $37811105.76)

= 7.1800%

Market values:

Debt = $64,837,229.52 + $37811105.76 = 102648335.28

Equity = 9,284,561 * 64.73 = 600989633.53

Hence:

WACC = 0.080333694 * 600989633.53 / (600989633.53+102648335.28) + 7.1800% * (1 - 34%) * 102648335.28 / (600989633.53+102648335.28)

= 0.0755

WACC = 0.0755